OpenStudy (anonymous):

y = (2/3)x - 4 Write an equation for the line passing through (5, 1) that is parallel to y = (2/3)x - 4.

jimthompson5910 (jim_thompson5910):

y = (2/3)x+b 1 = (2/3)(5) + b 1= 10/3 + b 1-10/3= b -7/3 = b b = -7/3 So the equation is \[\large y=\frac{2}{3}x-\frac{7}{3}\]